4.Consider the following information:

Evauate the foowing three projects, using the profitability index. Assume a cost of capita of 15 

Cash Flows Liquidate Recondition Replace

Project

A.Based on the given information, what would be the profitability index of each project respectivey?

Initial Cash Out fow –$100,000 –$500,000–$1,000,000

Year 3 cash inflow 75,000 250,000500,000

Year 2 cash inflow 60,000 200,000 500,000

Year 1 cash nfow 50,000 100,000 500,000

percent.

(Round your answer to the nearest four decimal points or use table value)


1
Expert's answer
2022-01-03T14:23:30-0500

Profitability Index (PI):

"PI=\frac{PV of future cash flows}{Initial Investment}"


First needs to determine present value (PV) of future cash flows for each project.


"PV(Liquidate )=\frac{50000}{(1.15)^1}+\frac{60000}{(1.15)^2}+\frac{75000}{(1.15)^3}=138161"

"PV(Recondition )=\frac{100000}{(1.15)^1}+\frac{200000}{(1.15)^2}+\frac{250000}{(1.15)^3}=402564"

"PV(Replace)=\frac{500000}{(1.15)^1}+\frac{500000}{(1.15)^2}+\frac{500000}{(1.15)^3}=1141613"


"PI(Liquidate )=\frac{138161}{100000}=1.3816"


"PI(Recondition )=\frac{402564}{500000}=0.8051"


"PI(Replace)=\frac{1141613}{1000000}=1.1416"


Profitability index (PI) of Liquidate project is highest among three projects.